# Multimodal Interpretation of Remote Sensing Images: Dynamic Resolution Input Strategy and Multi-scale Vision–Language Alignment Mechanism

### v1

@@ 1.1 Research significance of the subject
Remote sensing (RS) imagery, acquired via satellite-borne or airborne sensors, serves as a pivotal technology for capturing geospatial information of the Earth’s surface [ref1]. Endowed with inherent advantages including extensive coverage, high spatiotemporal resolution, and rich spectral characteristics, RS has become an indispensable tool across diverse domains such as environmental monitoring, urban planning, agricultural surveying, and disaster assessment [ref2, ref3]. With the continuous advancements in imaging and sensing technologies, RS data have evolved from traditional unimodal optical imagery to multimodal and multisource formats, encompassing synthetic aperture radar (SAR), light detection and ranging (LiDAR), multispectral, and hyperspectral imagery. This evolution has significantly enhanced the comprehensiveness and granularity of Earth observation, enabling more precise and holistic characterization of surface features [ref4, ref5, ref6].
@@
However, remote sensing (RS) imagery still confronts multiple technical bottlenecks in practical deployment: including complex background clutter, subtle inter-class discrepancies among land cover types, and significant time-varying fluctuations in illumination and meteorological conditions. Such inherent complexities have substantially constrained the performance of precise information extraction and high-level semantic interpretation tasks for RS data[ref7]. In this context, multimodal comprehensive understanding techniques offer a forward-looking technical pathway to address these challenges by fusing structural features, temporal dynamic patterns, and spatial correlations encapsulated in heterogeneous sensor data. This approach can effectively overcome the inherent limitations of unimodal RS analysis, enabling robust execution of high-level tasks (e.g., land cover classification, change detection, and scene semantic interpretation) and significantly improving the accuracy and reliability of intelligent RS imagery analysis[ref8, ref9].
@@
With the rapid advancement of deep learning and artificial intelligence technologies, a comprehensive multimodal understanding of RS imagery has emerged as a key frontier research direction in the contemporary RS domain[ref10]. On one aspect, constructing a unified multimodal data fusion framework enables both complementary information integration across heterogeneous data sources and effective suppression of uncertainties and noise perturbations associated with single-source data dependence, thus markedly improving the accuracy and robustness of land surface parameter retrieval and information extraction[ref11]. On the other aspect, to address practical challenges (e.g., strong spatiotemporal heterogeneity of RS data and scarcity of annotated samples), exploring emerging methodological paradigms (including self-supervised learning, cross-domain transfer learning, and graph neural networks) holds critical theoretical guidance and practical support value for enhancing model generalization and adaptability, as well as facilitating the large-scale deployment of multimodal RS technologies[ref12, ref13, ref14].
@@
In this context, this study presents two core innovations: a task-driven Dynamic Resolution Input Strategy (DRIS) and a Multi-scale Vision–language Alignment Mechanism (MS-VLAM). The former adaptively tunes the input image resolution based on task specifications and computational constraints, striking a balance between high analytical accuracy and computational efficiency. This attribute makes it particularly tailored to RS scenarios characterized by ultra-wide coverage and coexisting multi-scale features. The latter realizes cross-modal feature semantic alignment at the Object, Local-region, and Global levels, which markedly enhances the robustness and precision of cross-modal retrieval, object detection, and semantic interpretation for RS data[ref15]. This integrated design effectively alleviates the challenges of granularity inconsistency and semantic misalignment in multimodal fusion, thereby offering a novel paradigm for building an efficient and scalable RS multimodal understanding system.
@@
The advancement of multimodal comprehensive understanding technologies embodies not only substantial academic value but also broad prospects for practical and strategic applications. It facilitates the transition of RS data toward in-depth intelligent interpretation, driving technological progress in domains such as ecological conservation, smart agriculture, urban planning, and disaster response[ref16]. Specifically, in ecological and environmental conservation, multimodal fusion technologies enable more accurate monitoring of forest cover dynamics, wetland degradation processes, and water resource conditions (aligning with the framework’s ecological conservation application)[ref19]; in smart agriculture, they support dynamic crop growth monitoring and precise detection of pests and diseases[ref20]; in urban management, they assist in land use planning and surveillance of urban expansion (supporting land use planning and urban planning tasks)[ref21]; and in disaster response, they enhance rapid response and assessment capabilities for hazards including earthquakes, floods, and wildfires[ref22].
@@
Moreover, international cutting-edge research trends in remote sensing (RS) indicate that multimodal fusion-driven intelligent analysis is progressing toward three pivotal directions: large-scale spatiotemporal data processing (for global-scale Earth observation), cross-domain knowledge transfer (to mitigate data disparities across distinct RS scenarios), and multi-task joint learning (to unify heterogeneous downstream tasks)[ref57, ref58, ref59]. By harnessing integrated cloud-edge computing architectures, which integrate low-latency edge inference and large-capacity cloud storage, this paradigm enables real-time processing and on-demand updating of high-throughput RS data streams[ref60]. This capability directly overcomes the efficiency bottleneck of conventional RS information services, thereby further elevating the intelligence and automation of application-centric remote sensing service systems, e.g., those for regional ecological monitoring or emergency disaster response[ref61].
@@
In summary, research on multimodal comprehensive understanding of RS imagery is not only an inevitable requirement for the innovation and theoretical advancement of remote sensing technology but also a strategic initiative to address the demand for high-quality geospatial information services in the intelligent era. By providing theoretical guidance and technical support for constructing efficient and robust RS information processing systems, this research will contribute to the realization of Digital Earth and Smart Earth visions, while facilitating leapfrog development in geographic information science and related application domains[ref16, ref17].
@@ 1.2 Previous Methods and Shortcomings
As a core underpinning of visual-language understanding, the cross-modal alignment mechanism between vision and language has garnered extensive attention and in-depth exploration in recent years [ref18]. Existing methodologies primarily achieve cross-modal semantic matching between remote sensing (RS) imagery and text through feature representation learning and similarity measurement, yet they suffer from inherent limitations in two core technical aspects that hinder their performance in complex RS scenarios.
@@
On the one hand, conventional RS image feature extraction methodologies predominantly rely on fixed-resolution input strategies [ref23]. Specifically, most models normalize all input images to a pre-defined resolution to ensure consistent dimensions for batch processing and model training. Nevertheless, this paradigm exhibits notable drawbacks: high-resolution inputs preserve abundant fine-grained details but substantially augment computational overhead and GPU memory consumption, compromising the efficiency of both training and inference; in contrast, low-resolution inputs reduce computational costs but often lead to the loss of critical fine-scale visual information, impairing the model’s ability to perceive local details and thus undermining its performance in understanding complex RS scenes. Furthermore, the uniform fixed-resolution processing lacks flexibility to dynamically allocate computational resources according to the inherent complexity of RS images, ultimately restricting the model’s generalization across diverse RS visual interpretation tasks.
@@
On the other hand, existing RS cross-modal alignment approaches primarily rely on object-level or global-level strategies to establish correspondences between RS imagery and textual descriptions [ref24]. Early research focused predominantly on global alignment, which matches the holistic features of an entire RS image with the comprehensive semantic features of a complete text passage[ref25]. While this paradigm captures overall inter-modal semantic consistency, it overlooks inherent fine-grained semantic associations within both RS imagery and text , resulting in inadequate comprehension of specific ground objects or local scene details. Subsequent studies have attempted to integrate local feature alignment methods (e.g., region-based object detection bounding boxes or fragment-level feature representations) to enhance the matching accuracy between individual ground objects or local image regions and their corresponding textual elements[ref26]. However, most current approaches remain confined to object-level alignment, lacking systematic multi-scale semantic modeling capabilities that would enable simultaneous capture of semantic consistency across object-level, local-region-level, and global-level level. Moreover, object-level alignment often disregards the multi-layered semantic hierarchies inherent in both visual (RS imagery) and textual modalities, constraining the expressive power of cross-modal associations and rendering these methods ill-equipped for complex RS tasks demanding fine-grained semantic understanding and contextual reasoning, such as RS image-text retrieval and visual question answering.
@@
Beyond the aforementioned limitations, the majority of current RS cross-modal methods lack hierarchical structure in their overall feature representations [ref28], which hinders the effective capture of structured information regarding interrelationships among different ground objects, local regions, and their contextual associations. In complex RS scenes, ground objects typically exhibit specific spatial configurations, semantic correlations, and interactive relationships. These characteristics cannot be accurately depicted solely through flat and holistic feature descriptions. For advanced RS visual-language tasks (e.g., scene graph generation and multi-object interaction understanding), models require the ability to hierarchically recognize and represent distinct ground objects while capturing their positional, action, and causal relationships [ref27]. The single-level representation limitation of current methods restricts in-depth understanding and reasoning of inter-object relationships and semantic interactions in complex scenarios, failing to meet the requirements for high-level RS visual-language reasoning.
@@
Additionally, the trade-off between computational efficiency and alignment accuracy remains a prominent challenge for RS cross-modal alignment methods. Achieving finer-grained semantic alignment generally necessitates processing high-resolution image inputs and conducting complex multi-scale feature computations, leading to substantial consumption of computational resources and increased latency during both training and inference phases[ref62]. This constraint impairs the scalability and real-time performance of such methods in practical RS applications; particularly in large-scale RS datasets and real-time monitoring scenarios, the excessive computational burden becomes a critical bottleneck restricting the widespread deployment and adoption of cross-modal alignment models [ref29].
@@
In summary, existing RS vision-language alignment methods exhibit notable deficiencies across four interrelated aspects: inadequate fixed-resolution handling strategies that fail to balance efficiency and detail preservation [ref30]; insufficient multi-scale alignment paradigms that overlook inherent multi-layered semantic structures [ref31]; inadequate hierarchical modeling of complex ground objects and their interactive relationships [ref32]; and excessive computational resource demands that limit practical deployment [ref33]. To address these critical bottlenecks, it is essential to develop a cross-modal alignment framework that integrates dynamic resolution input processing with systematic multi-scale alignment. This framework specifically incorporates object-level, local-region-level, and global-level cross-modal visual-language alignment mechanisms, and it will serve as a key breakthrough for advancing high-performance visual-language understanding in RS.
@@ 2.1 Segmentation of Remote Sensing Images
Remote sensing (RS) image segmentation serves as a foundational step in intelligent RS scene interpretation. Its core objective is to partition complex RS scenes into semantically homogeneous regions and extract target object information, thereby supporting diverse downstream applications including urban planning, crop growth monitoring, and natural disaster assessment [ref34]. This section systematically reviews the evolutionary landscape of RS image segmentation methods, ranging from traditional techniques to state-of-the-art deep learning-based approaches, while clarifying key technical advancements, inherent limitations, and critical research gaps.
@@
Traditional RS image segmentation techniques were the earliest research focus, encompassing thresholding, edge detection, region growing and merging, graph-based algorithms (e.g., Graph Cut, Normalized Cut), and object-based image analysis (OBIA) [ref35]. These methods rely on handcrafted features such as pixel intensity, texture, and edge contours, and exhibit advantages of algorithmic simplicity, high computational efficiency, and strong interpretability. However, their performance degrades significantly when processing high-resolution RS imagery with complex backgrounds or multi-class objects, often leading to over-segmentation or under-segmentation artifacts. A fundamental limitation lies in their inability to effectively model contextual information and spatial structural relationships between objects.
@@
With the advancement of machine learning, RS image segmentation gradually shifted toward learning-based paradigms [ref36]. Representative methods include Support Vector Machines (SVM), Random Forests (RF), and clustering algorithms such as K-means and Fuzzy C-Means (FCM). These approaches reformulate segmentation as a pixel-level or object-level classification task, leveraging engineered spectral, texture, and shape features for discrimination. While they achieve improved segmentation accuracy in small-sample scenarios, these methods still depend heavily on labor-intensive manual feature engineering. Moreover, they lack the capability to automatically extract high-level semantic features, limiting their performance in capturing complex spatial contextual relationships within RS scenes.
@@
In recent years, deep learning-based methods have emerged as the dominant paradigm in RS image segmentation, driving substantial improvements in segmentation accuracy and spatial structure preservation. The Fully Convolutional Network (FCN) pioneered this field by replacing fully connected layers in traditional convolutional neural networks with convolutional layers, enabling end-to-end pixel-level segmentation and supporting input images of arbitrary sizes [ref37]. Despite this breakthrough, FCN suffers from severe spatial detail loss during successive downsampling operations, resulting in blurred segmentation boundaries that are unacceptable for high-precision RS applications.
@@
To address the limitations of FCN, U-Net and its variants have been widely adopted in RS image segmentation tasks. U-Net employs a symmetric encoder-decoder architecture, where the encoder extracts hierarchical features through downsampling and the decoder recovers spatial resolution via upsampling. Skip connections between corresponding encoder and decoder layers effectively fuse shallow fine-grained features with deep high-level semantic features, thus preserving boundary details and spatial information [ref38]. To adapt to the complex shapes and irregular distributions of ground objects in RS imagery, improved architectures such as Attention U-Net, ResU-Net, and Nested U-Net (U-Net++) have been proposed. These variants integrate attention mechanisms, residual connections, and dense connections to enhance the model’s capability in segmenting small objects and handling complex RS scenes [ref39].
@@
The DeepLab series represents another pivotal branch of deep learning-based RS segmentation methods, with DeepLabV3 and DeepLabV3+ being the most influential variants [ref40]. Their core innovations include Atrous Convolution and Atrous Spatial Pyramid Pooling (ASPP), which expand the receptive field of convolutional layers without increasing model parameters or computational costs. This design enables the model to capture multi-scale contextual information, making it suitable for segmenting RS targets of varying sizes [ref41]. In addition, DeepLabV3+ incorporates an encoder-decoder structure to further refine spatial details and adopts Conditional Random Fields (CRF) as a post-processing step to optimize boundary representations, significantly improving segmentation accuracy in complex RS scenes [ref42].
@@
With the growing demand for global context modeling in large-scale RS imagery, Transformer-based architectures and their hybrid designs have been increasingly introduced into RS image segmentation. Transformers leverage self-attention mechanisms to capture long-range spatial dependencies and global semantic relationships, providing strong interpretability for complex RS scenes [ref43]. Typical examples include Swin Transformer, which adopts a hierarchical structure and sliding window self-attention to balance global information capture and computational efficiency [ref44]; SegFormer, which combines a lightweight Transformer encoder with a simple decoder to achieve high-accuracy and efficient segmentation [ref45]; and TransUNet, which embeds Transformer modules into the U-Net architecture to integrate the local feature extraction capability of CNNs with the global modeling power of Transformers, enabling accurate segmentation in complex backgrounds [ref46].
@@
To further address the unique challenges of RS imagery, such as large target scale variations and blurred object boundaries, researchers have integrated a series of specialized enhancement techniques into deep learning models. Multi-scale feature fusion techniques including Feature Pyramid Networks (FPN), ASPP, and Pyramid Pooling Modules (PSPNet) are widely used to capture both local fine-grained details and global contextual information [ref47]. Attention mechanisms such as Squeeze-and-Excitation (SE) modules, Convolutional Block Attention Module (CBAM), and self-attention modules are employed to selectively enhance discriminative features, improving the model’s sensitivity to small objects and boundary regions [ref48]. For the problem of limited labeled RS data, weakly supervised segmentation, self-supervised pretraining, and domain adaptation methods have been explored to reduce reliance on large-scale annotated datasets and enhance cross-region generalization capability [ref49].
@@
Beyond single-modal and single-scale methods, multi-modal and multi-scale fusion approaches have been proposed to further boost segmentation performance for complex RS scenes. Multi-modal methods integrate complementary data sources such as optical imagery, Synthetic Aperture Radar (SAR) data, and LiDAR elevation data, performing fusion at the feature level, decision level, or joint embedding space to enhance the model’s discriminative ability in heterogeneous backgrounds [ref50]. Multi-scale methods adopt strategies including image pyramids, FPN, atrous convolutions, and ASPP to handle the scale diversity of RS targets, effectively addressing the low segmentation accuracy of small objects and large-scale scenes [ref51]. Additionally, Graph Neural Networks (GNNs) have been incorporated into RS segmentation by constructing spatial graph structures at the pixel or superpixel level, modeling spatial contextual relationships through node adjacency, and further improving segmentation completeness and accuracy [ref3].

### v2

@@ 1.1 Research significance of the subject
Remote sensing (RS) imagery, acquired via satellite-borne or airborne sensors, serves as a pivotal technology for capturing geospatial information of the Earth’s surface [39]. Endowed with inherent advantages including extensive coverage, high spatiotemporal resolution, and rich spectral characteristics, RS has become an indispensable tool across diverse domains such as environmental monitoring, urban planning, agricultural surveying, and disaster assessment [48, 31]. With the continuous advancements in imaging and sensing technologies, RS data have evolved from traditional unimodal optical imagery to multimodal and multisource formats, encompassing synthetic aperture radar (SAR), light detection and ranging (LiDAR), multispectral, and hyperspectral imagery. This evolution has significantly enhanced the comprehensiveness and granularity of Earth observation, enabling more precise and holistic characterization of surface features [58, 35, 40].
@@
However, remote sensing (RS) imagery still confronts multiple technical bottlenecks in practical deployment: including complex background clutter, subtle inter-class discrepancies among land cover types, and significant time-varying fluctuations in illumination and meteorological conditions. Such inherent complexities have substantially constrained the performance of precise information extraction and high-level semantic interpretation tasks for RS data[64]. In this context, multimodal comprehensive understanding techniques offer a forward-looking technical pathway to address these challenges by fusing structural features, temporal dynamic patterns, and spatial correlations encapsulated in heterogeneous sensor data. This approach can effectively overcome the inherent limitations of unimodal RS analysis, enabling robust execution of high-level tasks (e.g., land cover classification, change detection, and scene semantic interpretation) and significantly improving the accuracy and reliability of intelligent RS imagery analysis[34, 24].
@@
With the rapid advancement of deep learning and artificial intelligence technologies, a comprehensive multimodal understanding of RS imagery has emerged as a key frontier research direction in the contemporary RS domain[51]. On one aspect, constructing a unified multimodal data fusion framework enables both complementary information integration across heterogeneous data sources and effective suppression of uncertainties and noise perturbations associated with single-source data dependence, thus markedly improving the accuracy and robustness of land surface parameter retrieval and information extraction[65]. On the other aspect, to address practical challenges (e.g., strong spatiotemporal heterogeneity of RS data and scarcity of annotated samples), exploring emerging methodological paradigms (including self-supervised learning, cross-domain transfer learning, and graph neural networks) holds critical theoretical guidance and practical support value for enhancing model generalization and adaptability, as well as facilitating the large-scale deployment of multimodal RS technologies[22, 1, 56].
@@
In this context, this study presents two core innovations: a task-driven Dynamic Resolution Input Strategy (DRIS) and a Multi-scale Vision–language Alignment Mechanism (MS-VLAM). The former adaptively tunes the input image resolution based on task specifications and computational constraints, striking a balance between high analytical accuracy and computational efficiency. This attribute makes it particularly tailored to RS scenarios characterized by ultra-wide coverage and coexisting multi-scale features. The latter realizes cross-modal feature semantic alignment at the Object, Local-region, and Global levels, which markedly enhances the robustness and precision of cross-modal retrieval, object detection, and semantic interpretation for RS data[30]. This integrated design effectively alleviates the challenges of granularity inconsistency and semantic misalignment in multimodal fusion, thereby offering a novel paradigm for building an efficient and scalable RS multimodal understanding system.
@@
The advancement of multimodal comprehensive understanding technologies embodies not only substantial academic value but also broad prospects for practical and strategic applications. It facilitates the transition of RS data toward in-depth intelligent interpretation, driving technological progress in domains such as ecological conservation, smart agriculture, urban planning, and disaster response[17]. Specifically, in ecological and environmental conservation, multimodal fusion technologies enable more accurate monitoring of forest cover dynamics, wetland degradation processes, and water resource conditions (aligning with the framework’s ecological conservation application)[59]; in smart agriculture, they support dynamic crop growth monitoring and precise detection of pests and diseases[37]; in urban management, they assist in land use planning and surveillance of urban expansion (supporting land use planning and urban planning tasks)[4]; and in disaster response, they enhance rapid response and assessment capabilities for hazards including earthquakes, floods, and wildfires[20].
@@
Moreover, international cutting-edge research trends in remote sensing (RS) indicate that multimodal fusion-driven intelligent analysis is progressing toward three pivotal directions: large-scale spatiotemporal data processing (for global-scale Earth observation), cross-domain knowledge transfer (to mitigate data disparities across distinct RS scenarios), and multi-task joint learning (to unify heterogeneous downstream tasks)[18, 23, 44]. By harnessing integrated cloud-edge computing architectures, which integrate low-latency edge inference and large-capacity cloud storage, this paradigm enables real-time processing and on-demand updating of high-throughput RS data streams[43]. This capability directly overcomes the efficiency bottleneck of conventional RS information services, thereby further elevating the intelligence and automation of application-centric remote sensing service systems, e.g., those for regional ecological monitoring or emergency disaster response[53].
@@
In summary, research on multimodal comprehensive understanding of RS imagery is not only an inevitable requirement for the innovation and theoretical advancement of remote sensing technology but also a strategic initiative to address the demand for high-quality geospatial information services in the intelligent era. By providing theoretical guidance and technical support for constructing efficient and robust RS information processing systems, this research will contribute to the realization of Digital Earth and Smart Earth visions, while facilitating leapfrog development in geographic information science and related application domains[17, 10].
@@ 1.2 Previous Methods and Shortcomings
As a core underpinning of visual-language understanding, the cross-modal alignment mechanism between vision and language has garnered extensive attention and in-depth exploration in recent years [46]. Existing methodologies primarily achieve cross-modal semantic matching between remote sensing (RS) imagery and text through feature representation learning and similarity measurement, yet they suffer from inherent limitations in two core technical aspects that hinder their performance in complex RS scenarios.
@@
On the one hand, conventional RS image feature extraction methodologies predominantly rely on fixed-resolution input strategies [28]. Specifically, most models normalize all input images to a pre-defined resolution to ensure consistent dimensions for batch processing and model training. Nevertheless, this paradigm exhibits notable drawbacks: high-resolution inputs preserve abundant fine-grained details but substantially augment computational overhead and GPU memory consumption, compromising the efficiency of both training and inference; in contrast, low-resolution inputs reduce computational costs but often lead to the loss of critical fine-scale visual information, impairing the model’s ability to perceive local details and thus undermining its performance in understanding complex RS scenes. Furthermore, the uniform fixed-resolution processing lacks flexibility to dynamically allocate computational resources according to the inherent complexity of RS images, ultimately restricting the model’s generalization across diverse RS visual interpretation tasks.
@@
On the other hand, existing RS cross-modal alignment approaches primarily rely on object-level or global-level strategies to establish correspondences between RS imagery and textual descriptions [9]. Early research focused predominantly on global alignment, which matches the holistic features of an entire RS image with the comprehensive semantic features of a complete text passage[55]. While this paradigm captures overall inter-modal semantic consistency, it overlooks inherent fine-grained semantic associations within both RS imagery and text , resulting in inadequate comprehension of specific ground objects or local scene details. Subsequent studies have attempted to integrate local feature alignment methods (e.g., region-based object detection bounding boxes or fragment-level feature representations) to enhance the matching accuracy between individual ground objects or local image regions and their corresponding textual elements[62]. However, most current approaches remain confined to object-level alignment, lacking systematic multi-scale semantic modeling capabilities that would enable simultaneous capture of semantic consistency across object-level, local-region-level, and global-level level. Moreover, object-level alignment often disregards the multi-layered semantic hierarchies inherent in both visual (RS imagery) and textual modalities, constraining the expressive power of cross-modal associations and rendering these methods ill-equipped for complex RS tasks demanding fine-grained semantic understanding and contextual reasoning, such as RS image-text retrieval and visual question answering.
@@
Beyond the aforementioned limitations, the majority of current RS cross-modal methods lack hierarchical structure in their overall feature representations [60], which hinders the effective capture of structured information regarding interrelationships among different ground objects, local regions, and their contextual associations. In complex RS scenes, ground objects typically exhibit specific spatial configurations, semantic correlations, and interactive relationships. These characteristics cannot be accurately depicted solely through flat and holistic feature descriptions. For advanced RS visual-language tasks (e.g., scene graph generation and multi-object interaction understanding), models require the ability to hierarchically recognize and represent distinct ground objects while capturing their positional, action, and causal relationships [63]. The single-level representation limitation of current methods restricts in-depth understanding and reasoning of inter-object relationships and semantic interactions in complex scenarios, failing to meet the requirements for high-level RS visual-language reasoning.
@@
Additionally, the trade-off between computational efficiency and alignment accuracy remains a prominent challenge for RS cross-modal alignment methods. Achieving finer-grained semantic alignment generally necessitates processing high-resolution image inputs and conducting complex multi-scale feature computations, leading to substantial consumption of computational resources and increased latency during both training and inference phases[38]. This constraint impairs the scalability and real-time performance of such methods in practical RS applications; particularly in large-scale RS datasets and real-time monitoring scenarios, the excessive computational burden becomes a critical bottleneck restricting the widespread deployment and adoption of cross-modal alignment models [21].
@@
In summary, existing RS vision-language alignment methods exhibit notable deficiencies across four interrelated aspects: inadequate fixed-resolution handling strategies that fail to balance efficiency and detail preservation [49]; insufficient multi-scale alignment paradigms that overlook inherent multi-layered semantic structures [16]; inadequate hierarchical modeling of complex ground objects and their interactive relationships [33]; and excessive computational resource demands that limit practical deployment [19]. To address these critical bottlenecks, it is essential to develop a cross-modal alignment framework that integrates dynamic resolution input processing with systematic multi-scale alignment. This framework specifically incorporates object-level, local-region-level, and global-level cross-modal visual-language alignment mechanisms, and it will serve as a key breakthrough for advancing high-performance visual-language understanding in RS.
@@ 2.1 Segmentation of Remote Sensing Images
Remote sensing (RS) image segmentation serves as a foundational step in intelligent RS scene interpretation. Its core objective is to partition complex RS scenes into semantically homogeneous regions and extract target object information, thereby supporting diverse downstream applications including urban planning, crop growth monitoring, and natural disaster assessment [25]. This section systematically reviews the evolutionary landscape of RS image segmentation methods, ranging from traditional techniques to state-of-the-art deep learning-based approaches, while clarifying key technical advancements, inherent limitations, and critical research gaps.
@@
Traditional RS image segmentation techniques were the earliest research focus, encompassing thresholding, edge detection, region growing and merging, graph-based algorithms (e.g., Graph Cut, Normalized Cut), and object-based image analysis (OBIA) [45]. These methods rely on handcrafted features such as pixel intensity, texture, and edge contours, and exhibit advantages of algorithmic simplicity, high computational efficiency, and strong interpretability. However, their performance degrades significantly when processing high-resolution RS imagery with complex backgrounds or multi-class objects, often leading to over-segmentation or under-segmentation artifacts. A fundamental limitation lies in their inability to effectively model contextual information and spatial structural relationships between objects.
@@
With the advancement of machine learning, RS image segmentation gradually shifted toward learning-based paradigms [47]. Representative methods include Support Vector Machines (SVM), Random Forests (RF), and clustering algorithms such as K-means and Fuzzy C-Means (FCM). These approaches reformulate segmentation as a pixel-level or object-level classification task, leveraging engineered spectral, texture, and shape features for discrimination. While they achieve improved segmentation accuracy in small-sample scenarios, these methods still depend heavily on labor-intensive manual feature engineering. Moreover, they lack the capability to automatically extract high-level semantic features, limiting their performance in capturing complex spatial contextual relationships within RS scenes.
@@
In recent years, deep learning-based methods have emerged as the dominant paradigm in RS image segmentation, driving substantial improvements in segmentation accuracy and spatial structure preservation. The Fully Convolutional Network (FCN) pioneered this field by replacing fully connected layers in traditional convolutional neural networks with convolutional layers, enabling end-to-end pixel-level segmentation and supporting input images of arbitrary sizes [2]. Despite this breakthrough, FCN suffers from severe spatial detail loss during successive downsampling operations, resulting in blurred segmentation boundaries that are unacceptable for high-precision RS applications.
@@
To address the limitations of FCN, U-Net and its variants have been widely adopted in RS image segmentation tasks. U-Net employs a symmetric encoder-decoder architecture, where the encoder extracts hierarchical features through downsampling and the decoder recovers spatial resolution via upsampling. Skip connections between corresponding encoder and decoder layers effectively fuse shallow fine-grained features with deep high-level semantic features, thus preserving boundary details and spatial information [57]. To adapt to the complex shapes and irregular distributions of ground objects in RS imagery, improved architectures such as Attention U-Net, ResU-Net, and Nested U-Net (U-Net++) have been proposed. These variants integrate attention mechanisms, residual connections, and dense connections to enhance the model’s capability in segmenting small objects and handling complex RS scenes [3].
@@
The DeepLab series represents another pivotal branch of deep learning-based RS segmentation methods, with DeepLabV3 and DeepLabV3+ being the most influential variants [52]. Their core innovations include Atrous Convolution and Atrous Spatial Pyramid Pooling (ASPP), which expand the receptive field of convolutional layers without increasing model parameters or computational costs. This design enables the model to capture multi-scale contextual information, making it suitable for segmenting RS targets of varying sizes [32]. In addition, DeepLabV3+ incorporates an encoder-decoder structure to further refine spatial details and adopts Conditional Random Fields (CRF) as a post-processing step to optimize boundary representations, significantly improving segmentation accuracy in complex RS scenes [12].
@@
With the growing demand for global context modeling in large-scale RS imagery, Transformer-based architectures and their hybrid designs have been increasingly introduced into RS image segmentation. Transformers leverage self-attention mechanisms to capture long-range spatial dependencies and global semantic relationships, providing strong interpretability for complex RS scenes [27]. Typical examples include Swin Transformer, which adopts a hierarchical structure and sliding window self-attention to balance global information capture and computational efficiency [6]; SegFormer, which combines a lightweight Transformer encoder with a simple decoder to achieve high-accuracy and efficient segmentation [7]; and TransUNet, which embeds Transformer modules into the U-Net architecture to integrate the local feature extraction capability of CNNs with the global modeling power of Transformers, enabling accurate segmentation in complex backgrounds [5].
@@
To further address the unique challenges of RS imagery, such as large target scale variations and blurred object boundaries, researchers have integrated a series of specialized enhancement techniques into deep learning models. Multi-scale feature fusion techniques including Feature Pyramid Networks (FPN), ASPP, and Pyramid Pooling Modules (PSPNet) are widely used to capture both local fine-grained details and global contextual information [41]. Attention mechanisms such as Squeeze-and-Excitation (SE) modules, Convolutional Block Attention Module (CBAM), and self-attention modules are employed to selectively enhance discriminative features, improving the model’s sensitivity to small objects and boundary regions [13]. For the problem of limited labeled RS data, weakly supervised segmentation, self-supervised pretraining, and domain adaptation methods have been explored to reduce reliance on large-scale annotated datasets and enhance cross-region generalization capability [36].
@@
Beyond single-modal and single-scale methods, multi-modal and multi-scale fusion approaches have been proposed to further boost segmentation performance for complex RS scenes. Multi-modal methods integrate complementary data sources such as optical imagery, Synthetic Aperture Radar (SAR) data, and LiDAR elevation data, performing fusion at the feature level, decision level, or joint embedding space to enhance the model’s discriminative ability in heterogeneous backgrounds [26]. Multi-scale methods adopt strategies including image pyramids, FPN, atrous convolutions, and ASPP to handle the scale diversity of RS targets, effectively addressing the low segmentation accuracy of small objects and large-scale scenes [50]. Additionally, Graph Neural Networks (GNNs) have been incorporated into RS segmentation by constructing spatial graph structures at the pixel or superpixel level, modeling spatial contextual relationships through node adjacency, and further improving segmentation completeness and accuracy [31].